# Thermodynamic uncertainty relations for relativistic quantum thermal machines

### v1

@@ I Introduction
Advancements in nanoscale device fabrication and the coherent control of atoms have paved the way for the realization of thermal machines operating in the quantum regime, enabling the investigation into the thermodynamic behavior of quantum systems and the role of quantum information in thermodynamic processes [2, 3, 4, 5, 6, 7]. At the microscopic scale, stochastic fluctuations of thermodynamic variables and irreversible entropy production become inherently significant, often affecting the performance of quantum thermal machines. Consequently, elucidating the fundamental principles that govern nonequilibrium thermodynamic processes is of crucial importance. In this context, various symmetry relations, known as fluctuation theorems [8, 9, 10, 11, 12, 13, 14], have been formulated that characterize the probability distributions of stochastically fluctuating quantities. They provide an extension of the second law of thermodynamics to the framework of stochastic thermodynamics, quantifying the occurrence of negative fluctuations of entropy.
@@
An UDW detector moving at a constant relativistic speed experiences a frequency-dependent effective temperature, which can be either hotter or colder than the ambient temperature of the thermal field bath. Building on this observation, we previously introduced a quantum SWAP heat engine (a two-stroke engine) [58, 59, 60, 61, 62], where the two-qubit working medium is modeled by inertially moving UDW detectors, each locally coupled to a thermal field bath at a different temperature. Here, we derive generalized efficiency bounds that govern the performance of the thermal machine operating either as a heat engine or as a refrigerator. These bounds arise from the second law of stochastic thermodynamics. We identify regimes in which relativistic motion leads to stronger violations of the classical TUR bound, previously observed in the quantum regime.
@@ III Otto cycle with moving qubits
We consider a two-stroke quantum SWAP thermal machine [59, 60, 61, 62], consisting of two qubits labeled $A$ and $B$, with respective transition frequencies $\omega_{A}$ and $\omega_{B}$. Each qubit is described by the Hamiltonian $H_{i}=\omega_{i}\sigma_{z}^{i}/2$, with $i\in\{A,B\}$. Initially, qubit $A$ is brought into thermal equilibrium with a hot reservoir at temperature $T_{A}$, while qubit $B$ with a cold reservoir at $T_{B}$ (that is, $T_{A}>T_{B}$). The initial state of the two-qubit system is
@@
We now consider that the two-qubit working medium of the engine is represented by two UDW detectors moving at constant speeds $\upsilon_{A}$ and $\upsilon_{B}$. During the first stroke of the cycle, the moving qubits $A$ and $B$ interact with hot and cold thermal field baths at temperatures $T_{A}$ and $T_{B}$ (with $T_{A}>T_{B}$), respectively. Due to their motion, each qubit experiences the effective temperatures denoted by $T^{\text{eff}}_{A}$ and $T^{\text{eff}}_{B}$. Once both qubits reach their respective steady states $\rho_{D}$, with $D\in\{A,B\}$, the engine proceeds through the subsequent strokes of the cycle as in the static configuration.
@@ IV Thermodynamic uncertainty relations
In Fig. 1, we present the fluctuations in heat exchanged with the hot bath, $\text{Var}[Q_{H}]$, for the case where either qubit A or B moves with constant relativistic velocity. The results are compared against the derived bound (22), the classical TUR bound (1), as well as the generalized TUR bound introduced in [61]. The latter is given by
@@
where $f(x)=\text{csch}^{2}(g(x/2))$, with $\text{csch}(x)\equiv 1/\sinh x$, and $g(x)$ denoting the function inverse of $x\tanh(x)$. In the limit $\expectationvalue{\Sigma}\ll 1$, the classical TUR (1) is recovered from Eq. (23). We observe that the TUR bound can be violated for small values of the transition frequency ratio $\omega_{B}/\omega_{A}$, whereas both the bound (22) and the generalized bound (23) remain valid.
@@
for different qubit velocities $\upsilon_{A}$ and $\upsilon_{B}$, as a function of $\omega_{B}\beta_{B}$ with $\omega_{A}\beta_{A}=0.5$ fixed, and as a function of $\omega_{A}\beta_{A}$ with $\omega_{B}\beta_{B}=0.5$ fixed. This illustrates how the TUR bound is modified when the qubits are in motion compared to the rest-frame scenario ($\upsilon_{A}\!=\!\upsilon_{B}\!=\!0$). We observe that relativistic motion can lead to stronger violations of the bound already observed in the quantum regime. In particular, motion through the hot bath increases the violation when the machine operates as a refrigerator, whereas motion through the cold bath leads to a larger violation when the machine functions as a heat engine. These results suggest that relativistic motion may be harnessed as a resource to operate the machine beyond classical bounds.
@@ V Generalized performance bounds
Now, we present performance bounds when the machine is functioning as a heat engine and, then a refrigerator. The efficiency, $\eta$, of the heat engine mode of operation, defined as the ratio of the total work output to the absorbed heat, is given by
@@
The uncertainty relation (22) then yields the following trade-off between the engine’s efficiency and mean power output
@@
This implies that reaching efficiencies close to the generalized Carnot limit requires either a reduction in power output or an increase in the second moment of the work distribution. In our previous work [58], we optimized the performance of the heat engine operating at finite power by evaluating its efficiency at maximum power, demonstrating that relativistic motion through the thermal field baths can yield an efficiency that exceeds the standard Carnot efficiency defined with respect to the rest-frame temperatures (see also [80]).

### v2

@@ I Introduction
Advancements in nanoscale device fabrication and the coherent control of atoms have paved the way for the realization of thermal machines operating in the quantum regime, enabling the investigation into the thermodynamic behavior of quantum systems and the role of quantum information in thermodynamic processes [2, 3, 4, 5, 6, 7]. At the microscopic scale, stochastic fluctuations of thermodynamic variables and irreversible entropy production become inherently significant, often affecting the performance of quantum thermal machines. Consequently, elucidating the fundamental principles that govern nonequilibrium thermodynamic processes is of crucial importance. In this context, various symmetry relations, known as fluctuation theorems [8, 9, 10, 11, 12, 13, 14], have been formulated to characterize the probability distributions of stochastically fluctuating quantities. They provide an extension of the second law of thermodynamics to the framework of stochastic thermodynamics, quantifying the occurrence of negative fluctuations of entropy.
@@
An UDW detector moving at a constant relativistic speed experiences a frequency-dependent effective temperature, which can be either hotter or colder than the ambient temperature of the thermal field bath. Building on this observation, we previously introduced a quantum SWAP heat engine (a two-stroke engine) [58, 59, 60, 61, 62, 63], where the two-qubit working medium is modeled by inertially moving UDW detectors, each locally coupled to a thermal field bath at a different temperature. Here, we derive generalized efficiency bounds that govern the performance of the thermal machine operating either as a heat engine or as a refrigerator. These bounds arise from the second law of stochastic thermodynamics. We identify regimes in which relativistic motion leads to stronger violations of the classical TUR bound, previously observed in the quantum regime.
@@ III Otto cycle with moving qubits
We consider a two-stroke quantum SWAP thermal machine [59, 60, 61, 62, 63], consisting of two qubits labeled $A$ and $B$, with respective transition frequencies $\omega_{A}$ and $\omega_{B}$. Each qubit is described by the Hamiltonian $H_{i}=\omega_{i}\sigma_{z}^{i}/2$, with $i\in\{A,B\}$. Initially, qubit $A$ is brought into thermal equilibrium with a hot reservoir at temperature $T_{A}$, while qubit $B$ with a cold reservoir at $T_{B}$ (that is, $T_{A}>T_{B}$). The initial state of the two-qubit system is
@@
We now consider that the two-qubit working medium of the engine is represented by two UDW detectors moving at constant speeds $\upsilon_{A}$ and $\upsilon_{B}$. As in the standard UDW detector framework (e.g., for uniformly accelerated detectors), the relativistic motion of the detectors is assumed to be externally sustained and their trajectories fixed, without explicitly modelling the associated energetic cost or backreaction. During the first stroke of the cycle, the moving qubits $A$ and $B$ interact with hot and cold thermal field baths at temperatures $T_{A}$ and $T_{B}$ (with $T_{A}>T_{B}$), respectively. Due to their motion, each qubit experiences the effective temperatures denoted by $T^{\text{eff}}_{A}$ and $T^{\text{eff}}_{B}$. Once both qubits reach their respective steady states $\rho_{D}$, with $D\in\{A,B\}$, the engine proceeds through the subsequent strokes of the cycle as in the static configuration.
@@ IV Thermodynamic uncertainty relations
In Fig. 1, we present the fluctuations in heat exchanged with the hot bath, $\text{Var}[Q_{H}]$, for the case where either qubit A or B moves with constant relativistic velocity. The results are compared against the derived bound Eq. (22), the classical TUR bound Eq. (1), as well as the generalized TUR bound introduced in [61]. The latter is given by
@@
where $f(x)=\text{csch}^{2}(g(x/2))$, with $\text{csch}(x)\equiv 1/\sinh x$, and $g(x)$ denoting the function inverse of $x\tanh(x)$. A Taylor expansion of $f(\expectationvalue{\Sigma})$ yields $f(\expectationvalue{\Sigma})\simeq 2/\expectationvalue{\Sigma}-2/3$, and thus for $\expectationvalue{\Sigma}\ll 1$ the bound in Eq. (1) is recovered [61]. We observe that the classical TUR bound can be violated for small values of the transition frequency ratio $\omega_{B}/\omega_{A}$, whereas both the bound Eq. (22) and the generalized bound Eq. (23) remain valid.
@@
for different qubit velocities $\upsilon_{A}$ and $\upsilon_{B}$, as a function of $\omega_{B}\beta_{B}$ with $\omega_{A}\beta_{A}=0.5$ fixed, and as a function of $\omega_{A}\beta_{A}$ with $\omega_{B}\beta_{B}=0.5$ fixed. This illustrates how the TUR bound is modified when the qubits are in motion compared to the rest-frame scenario ($\upsilon_{A}\!=\!\upsilon_{B}\!=\!0$). We observe that relativistic motion can lead to stronger violations of the bound already observed in the quantum regime. In particular, motion through the hot bath increases the violation when the machine operates as a refrigerator, whereas motion through the cold bath leads to a larger violation when the machine functions as a heat engine. These results suggest that relativistic motion may be harnessed as a resource to operate the machine beyond classical bounds (see also Refs. [58, 81, 82]).
@@ V Generalized performance bounds
Now, we present performance bounds when the machine is functioning as a heat engine and, then as a refrigerator. The efficiency, $\eta$, of the heat engine mode of operation, defined as the ratio of the total work output to the absorbed heat, is given by
@@
The uncertainty relation Eq. (22) then yields the following trade-off between the engine’s efficiency and mean power output
@@
This implies that reaching efficiencies close to the generalized Carnot limit requires either a reduction in power output or an increase in the second moment of the work distribution. In our previous work [58], we optimized the performance of the heat engine operating at finite power by evaluating its efficiency at maximum power, demonstrating that relativistic motion through the thermal field baths can yield an efficiency that exceeds the standard Carnot efficiency defined with respect to the rest-frame temperatures.